6.f(x) = x–3/x+2 6.g(x) = 3x2/x2 + x - 6

x-intercepts
y-intercepts
Vertical asymptotes
Horizontal asymptotes
Domain of the function

Answer by stanbon(75887)

6.f(x) = x–3/x+2
---------------
x-intercepts:
Let y=0; solve for x as follows:
Y=0 when x-3=0 or x=3
x-intercept: (3,0)
-------------
y-intercepts
Let x=0 then y=-3/2
y-intercept: (0,-3/2)
--------------
Vertical asymptotes when the denominator is zero, i.e. when x=-2
-----------
Horizontal asymptotes at y=1/1=1; the 1/1 comes from the coefficients of x)
-------------
Domain of the function
All Real x except x=-2
---------------
6.g(x) = 3x^2/x^2 + x - 6
x-intercepts
Let y=0; y is zero is 3x^2=0; x-int at x=0
-----------
y-intercepts
Let x=0, then y=0
--------------
Vertical asymptotes when x^2+x-6=0
when (x+3)(x-2)=0, when x=-3 or x=2.
----------------
Horizontal asymptotes
Coefficents of x^2 are 3/2 ; Asymptote at y=3/2
------------------
Domain of the function
All Real numbers except x=-3 and x=2
---------------
Cheers,
Stan H.